OpenStudy (anonymous):

Evaluate cos2π/3−arctan1

OpenStudy (anonymous):

I need help on this maths

OpenStudy (hba):

\[\cos \frac{ 2 \pi }{ 3 }-\arctan(1)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

arctan(1) = pi/4 +npi where n is any integer

OpenStudy (anonymous):

u know the value of cos(2pi/3)?

OpenStudy (anonymous):

no idea

OpenStudy (hba):

If you wanna know its cos (120)

OpenStudy (anonymous):

cos(120) = cos(180-60) = -cos60

OpenStudy (anonymous):

Learning the values of sin and cosin for the unit circle would be ideal. http://www.mathsisfun.com/geometry/unit-circle.html

OpenStudy (anonymous):

u know cos60 ?

OpenStudy (anonymous):

they provided me with some options

OpenStudy (anonymous):

cos2π/3−arctan1 -1/2-pi/4 -n* pi

OpenStudy (anonymous):

what are ur options?

OpenStudy (anonymous):

this is one example 1−√3/2

OpenStudy (anonymous):

3√−1/2 another option